OpenStudy (anonymous):

hey I need to find out whether a few differential equations are separable, first order linear, both, or neither. one of the equations is y'+4y^2=0

OpenStudy (watchmath):

non linear since we have a y^2 there and it is separable since we can rewrite as dy/(-4y^2)=dt

OpenStudy (anonymous):

ok I understand the linear equation part, but not so much the separable part.

OpenStudy (watchmath):

what is your definition for separable differential eq.?

OpenStudy (anonymous):

an equations that has y'=f(x)g(y)

OpenStudy (watchmath):

ok, then y'=(-4)y^2 and you can take f(x)=-4 and g(y)=y^2

OpenStudy (anonymous):

ok thanks, and for first order linear just look for highest power of y?

OpenStudy (anonymous):

some of the equations have x's

OpenStudy (watchmath):

not the power of y but what is the highest derivative that you have for example y'''+xy'-x^3=e^x is linear of order 3

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

would y''+x^2y+z=0 be separable?